Find the angular displacement of a particle in 6 sec on a circle with a angular velocity 2π/3

Grade:12

2 Answers

Vikas TU
7 years ago
Angular displacement of a particle in 6 sec on a circle with a angular velocity 2π/3 is:
thetha = w*t = > 2pi*6/3 = > 4pi rad.
Ashmak Moon
7 years ago
There is a difference between angular distance and angular displacement, angular displacement is a vector quantity with magnitude equal to the shortest distance between initial and final positions of a particle and the direction is from initial to final position. in our case, the distance moved is complete two circomference '4pi’ and hence displacement is a null vector (a vector with zero magnitude and undetermined (not necessarily) direction. In our example, the direction is same as the direction of the angular velocity,i.e., clockwise or anticlockwise.